OpenStudy (anonymous):

Find the distance between the points (-3, 2) and (4, -5).

OpenStudy (anonymous):

use distance formula and you should get an answer of 9.899494937 so approx 10 rounded up

OpenStudy (anonymous):

thank you so much for the help !

OpenStudy (anonymous):

On an coordination plane P1=(x,y) P2=(z,t) Distance between P1 and P2: \[\sqrt{(t-y)^{2}+(z-x)^{2}}\] So , \[\sqrt{(-5-2)^{2}+(4+3)^{2}}=7\sqrt{2}\]